The Newman Corporation is interested in examining its cash conversion cycle. Suppose a Mennen manager has assembled the following data for your use: The inventory turnover is 14 times a year, the receivables turnover is 10 times a year, and the payables turnover is 12 times a year. What is the cash conversion cycle (CCC)? (Enter your answer as a decimal accurate to four decimal places)
Answer #1

Inventory period=365/inventory turnover

=365/14=26.0714286 days

Days in receivables=365/receivable turnover

=365/10=36.5 days

Days in payables=365/payable turnover

=365/12=30.4166667 days

Cash conversion cycle=Inventory period+Days in receivables-Days in payables

=26.0714286+36.5-30.4166667

=32.1548 days(Approx).